OpenStudy (anonymous):

8. The cost of four scarves and six hats is $52.00. The cost of two hats is $1.00 more than the cost of one scarf. What is the cost of one scarf? (1 point) $4.00 $5.00 $6.00 $7.0

TheSmartOne (thesmartone):

Ok, so we need to make a system of equations to solve this. :) What variable do you want to give to scarves and what variable do you want to give to hats?

TheSmartOne (thesmartone):

@CamoGreen

OpenStudy (anonymous):

Sorry, um s for scarves and h for hats

OpenStudy (anonymous):

@TheSmartOne

OpenStudy (anonymous):

C

OpenStudy (anonymous):

How did you get that answer?

OpenStudy (anonymous):

@xDTBxBashy2004

OpenStudy (anonymous):

Okay first equation 4s+6h=52 @TheSmartOne

TheSmartOne (thesmartone):

Correct @CamoGreen

OpenStudy (anonymous):

Okay then the next equation would be 2h=s+1 Then you would subtarct s and get 2h-s=1. Right @TheSmartOne

OpenStudy (anonymous):

So.... 4s+6h=52 > switch to... 6h+4s=52 2h-s=1

TheSmartOne (thesmartone):

here's a better idea

OpenStudy (anonymous):

k im all ears or eyes

TheSmartOne (thesmartone):

why don't we make the equation from: 2h = s+1 to: s = 2h - 1

OpenStudy (anonymous):

ok

TheSmartOne (thesmartone):

then we have the 2 equations 6h + 4s = 52 s = 2h - 1 we can then plug in equation #2 into equation #1

OpenStudy (anonymous):

O! Substitution instead of elimination. I like that better. Lol

OpenStudy (anonymous):

Ya I'll go that way

TheSmartOne (thesmartone):

good :)

OpenStudy (anonymous):

6h+4(2h-1)=52 6h+8h-4=52 14h-4=52 14h-4+4=52+4 14h=56 14h/14=56/14 h=4

OpenStudy (anonymous):

Then plug it in to one of the equations. 6(4)+4s=52 > 24+4s=52 > 24+4s-24=52-24> 4s=28 > 4s/4=28/4 > s=7. Right? @TheSmartOne

OpenStudy (anonymous):

So h=4 and s=7. 6h+4s=52 > 6(4)+4(7)=52 > 24+28=52 > 52=52 --Correct!

TheSmartOne (thesmartone):

Great job! :D

OpenStudy (anonymous):

Thank you and your brilliant mind! I never would have thought to use substitution! Thank ya!

OpenStudy (anonymous):

Thanks!

TheSmartOne (thesmartone):

Anytime! And great job once again! :D

OpenStudy (anonymous):

Haha thanks.